Dublin University Rifle Club

Version Date : May 9, 2001 Filename : Status :

Dublin University Rifle Club Chief Range Officer’s Report

Mark Dennehy Chief Range Officer

May 9, 2001

1 Membership and Activity levels

Membership of the club at the time of the AGM stands at 335 members. of which some 270 joined in October, 251 of which joined during Fresher’s Week. This represents an increase of over a hundred members from last year, and a huge improvement in the Fresher’s Week numbers, but also the third highest membership figure the club has had in the past twenty-five years, and it makes us the fourth-largest sports club in College. Looking at the membership figures for the past two decades or so shows that this represents an upward trend (see Figure 1) despite a certain amount of annual fluctuation, as well as a slump in membership over the past few years caused by the departure of a large group of our core shooters at the end of 1994.

Of these 335 members, 110 have shot at least once, giving an active membership of 32.8%. A closer look at the activity statistics shows that our most active term was Hillary term, as per usual, but also that there were relatively few people that only shot once - some 5 or 6 out of 110. This indicates that if we can get people to come down and try at least once, that they tend to return to try again - something to note for next year.

Looking at the demographics, we see that we have approximately a 2:1 ration of men to women in the club - this has been the average for some time now, but given that the lady shooters tend to be better than the gentlemen, this is a situation we would like to see changed to a 1:1 ratio.

The most well-represented faculty was Engineering, with 50 engineering students as members, and a further 30 in various management science courses, and another 32 in computer science and ICT. Arts was the next faculty, with 50 members, followed by Science with 36 and Medecine with 29. The other members fall into the “other” category, including Alumni and Staff members. In general, most of our members were Junior Freshmen (a total of 147) with the Senior Freshmen, Sophisters and Postgrads all roughly equal on approximately

1 Membership and Activity levels Figure 1: DURC membership trend since 1970

2 Safety

30-40 members each. Our ages range from 17 (when joining the club) to 55 years old - though the majority of shooters are currently 18 or 19, with the next largest group being the over-40s - a rather unique demographic amoung TCD’s sports clubs!.

This year’s has see us use a total of 19289 rounds of .22 ammunition, of which approximately 10,000 were shot on normal training nights and the rest in competition. This is up significantly on previous years, but we still have a ways to go to meet the previous record levels of 40,000 rounds per year (which were set in 1987-89)

Looking at the top ten of the ladder (the standard for judgeing our performance for the year) shows an average of 92.89 : not as high as our halycon days, but very respectable indeed. Looking at the last 12 cards shot shows a better picture of our current standards though : 93.38 as an average, with the top 8 having an average over 90. Also to note are the three possibles shot this year : congradulations to Ray Reilly, John Keeney and me. ☺

2 Safety

Despite an incident John has mentioned, this has been yet another incident-free year, with no injuries apart from the odd bruised head thanks to the ceiling of the Range. This means that for the past 39 years we have operated, we have had a total of zero injuries.

3 Courses

This year has seen the introduction of two new courses run by the National Target Shooting Association (N.T.S.A.) and the National Rifle and Pistol Association of Ireland (N.R.P.A.I.), which were lobbied for by Dublin University Rifle Club (D.U.R.C.) and University College Dublin Rifle Club (U.C.D.R.C.).

A Coaching course was held in U.C.D.R.C. in early September, with UK National Coaching Director Dave Froggett as instructor. D.U.R.C. sent four Range Officers to this course, which proved to be extremely useful both for these Officers’ personal shooting and for their ability to coach the Club’s Ordinary Members. A second course was scheduled to focus on smallbore shooting, but has been postponed due to the Foot and Mouth crisis. It will take place later this year.

A course for Range Officers has been lobbied for for several years by both D.U.R.C. and U.C.D.R.C., and more recently by WTSC. The N.R.P.A.I. had run such a course in the past (in the eighties), but had not done so recently, until this year. D.U.R.C. had been planning on sending all Range Officers to this course for certification, but confusion over the time and date of the course resulted in a very poor turnout on the day. The course will be rerun later this year and it is hoped that all Range Officers will become nationally certified before the end of the next Academic Year. As it is planned for this course to be an ongoing one run by the National Governing Body, it is hoped that all future Range Officers will be sent on it before they make the transition from Trainee to Full Range Officer.

4 Range Officers

This year has seen the departure of several Range Officers, including some of the old guard. Doctor Ian Marnane graduated early in 2000, and left soon thereafter, though not before doing a night or two of RO Duty, for old times sake. Leaving with him was long-standing member Andrea Lazenby-Simpson, who graduated last year but volunteered to assist in extremis. Mark Treanor also graduated and decided that it was a better idea to earn money in the real world than become a postgraduate! The CRO at the start of the year, Willie O’Connor, had to leave due to family and work committments, at which time I took over as CRO – filling his shoes proved a difficult task. Eric Farrell also had to leave, due to the workload his studies were imposing - hopefully he may decide to return at a future date.

5 Thanks

This loss of ROs does happen from year to year, but this year proved to be especially heavy and left us rather shorthanded - a larger number of TROs than usual were recruited as a result. The new Range Officers this year were David Gallagher and Shane Kelly (who had been Trainee Range Officers the year before), Peter Byrne, Conor McDermottroe, Liam Murray and Stephen Watterson. The Committee hereby offically welcomes them to a long and happy period of servitude in the name of the Club ☺.

5 Thanks

The first thanks of this report belongs squarely with the Range Officers who served this year, especially during Michaelmas and Hillary Terms; the level of work that they accepted and the relative lack of complaints were both quite exceptional The Club owes them a debt of gratitude – had they not stood RO Duty every week (or twice a week in some cases), the Club could not have continued through Michaelmas Term, let alone all the way to Trinity term.

Next year’s Range Officers won’t have to work quite so hard – but they are no less important to the operation of the Club and I’d like to thank them for volunteering.

A few specific mentions shoudl be made for :

• Keith White, for his occasional small piece of DIY on the range (ie, rebuilding it every so often), as well as for having written SCORP and KADA, the programs that we use to score competitions and track ladder scores throughout the year.

• Our “transport officers”, who volunteered to drive people and equipment all over the country to competitions: Ray Reilly (who was the sole transport officer last year), Mark Dennehy (yay me!), Jason Sibley, Alan Smyth (despite a recent reduction in the amount of gear he can carry), and our newest TO, Peter Byrne, showing how you can fit a big guy in a tiny little kiddies car ☺

• The Range Officers who cleaned and painted the Range last October so that it would be nice and clean for this year - it was a lot of hard work, taking nearly a full fortnight to complete under somewhat unpleasant conditions, especially when cleaning our the targets’ bullet traps. They were : Keith White, Mark Dennehy (yay me!), Ben Jones and Alan Smyth.

• All the Range Officers who worked on the Freshers Week stand - we took in 279 memebers this year, which was a huge improvement over previous years (and not just last year, when noone showed up).

• John Keeney for his level-headed leadership throughout the year, but more for his lack of delegation.

Thanks are also due to Tim Cooper and his Office; to the Chief Steward Pat Moray; and to the Security guards, who were their usual friendly and helpful selves throughout the year.

Thanks go to the Gardai, specifically Superintendent Dan Quill and Sergeant John Rafter, who wrote up our Removal Orders and other offical paperwork throughout the year without a minimum of fuss: and also to the RUC Firearms division, specifically Hillary Orr, who put through more than one rushed application for permits for competitions in Northern Ireland: and to Pam Temple of the Isle of Man Police Firearms Office, who managed to get two firearms permits sorted out with a total of fifteen minutes remaining before the Office closed, allowing me to go the Isle of Man easter shoot.

Both DUCAC and the College’s Sports Department deserve thanks, especially the Administrative Officers, Cathy and Geraldine, who have been of enormous assistance whenever it was humanly possible.

And finally, the last word of thanks is simply to all of the Club Members who showed up and shot during the year, especially if you shot at a competition. This is what keeps the Club alive from year to year, and I hope you will all be doing so again next year.

Acronyms

Acronyms

D.U.R.C. Dublin University Rifle Club U.C.D.R.C. University College Dublin Rifle Club N.T.S.A. National Target Shooting Association N.R.P.A.I. National Rifle and Pistol Association of Ireland